archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    a fiber cut off in Marseille the site MEN13 is inaccessible The intervention on the fiber is currently in progress Date Sunday 06 April 2014 02 05AM Reason for closing Done Comments 2 Comment by OVH Saturday 05 April 2014

    Original URL path: http://status.ovh.es/?do=details&id=6662 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    Courbeboie on the night of the 25th 26th in order to correct a bug on the management card Date Wednesday 26 March 2014 09 48AM Reason for closing Done Comments 4 Comment by OVH Wednesday 26 March 2014 09 47AM We will start the maintenance on the DSLAM Comment by OVH Wednesday 26 March 2014 09 47AM We are saving the configuration Comment by OVH Wednesday 26 March 2014 09

    Original URL path: http://status.ovh.es/?do=details&id=6594 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Details We are having a fiber cutoff between OVH and the NRA Hellemmes The NRA LQN59 LZI59 ASC59 TRI59 and HEL59 are no longer secured There s no impact on the traffic the DWDM security did its work and allows to keep the internet access Date Saturday 22 March 2014 22 22PM Reason for closing Done Comments 1 Comment by OVH Saturday 22 March 2014 22 22PM The security is

    Original URL path: http://status.ovh.es/?do=details&id=6573 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details We will be carrying out maintenance tonight on NRA WAS59 and SAN59 A few minutes downtime is expected Date Thursday 20 March 2014 01 11AM Reason for closing Done Comments 2 Comment by OVH Thursday 20 March

    Original URL path: http://status.ovh.es/?do=details&id=6554 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    PPP sessions for this collect are affected Date Friday 14 March 2014 09 25AM Reason for closing Done Comments 3 Comment by OVH Friday 14 March 2014 09 23AM We are no longer receiving the optical signal on the interco port with Axione Comment by OVH Friday 14 March 2014 09 24AM The Axione operator is migrating moving its equipments on the TH2 site where we have our collect port

    Original URL path: http://status.ovh.es/?do=details&id=6534 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Task Type Modernization Category Backend Core Status CLOSED Percent Complete Details We have some feedback on a malfunctioning of IP 29 blocks We are currently modifying the configuration in order to correct the default no service downtime is expected Date

    Original URL path: http://status.ovh.es/?do=details&id=6531 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Done Comments 6 Comment by OVH Wednesday 12 March 2014 09 37AM The teams are intervening on the pro75 site which is currently unreachable Comment by OVH Wednesday 12 March 2014 10 50AM The pro75 DSLAM is accessible again The intervention continuse to restore the securisation Comment by OVH Wednesday 12 March 2014 13 04PM The teams are now intervening on the gut75 fibre optics The DSLAM has been isolated

    Original URL path: http://status.ovh.es/?do=details&id=6523 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Details After a testing phase on the TG788v2s in particular a configuration update of the modems TG784 TG788 and TG789 will be deployed tonight This will improvement QoS management on the modem especially OVH VoIP The update is being carried out without service downtime and only on the modems in router mode Date Tuesday 11 March 2014 21 13PM Reason for closing Done Comments 2 Comment by OVH Tuesday 11

    Original URL path: http://status.ovh.es/?do=details&id=6518 (2015-10-15)
    Open archived version from archive



  •